# TV-TREES: Multimodal Entailment Trees for Neuro-Symbolic Video Reasoning

### v3

@@ 7 Conclusion
We introduce the first neuro-symbolic entailment tree generator for multimodal content to improve the robustness, reliability, interpretability, and scalability of video-language understanding systems. We focus on the application of narrative-driven VideoQA and show that our approach achieves state-of-the-art results on the zero-shot TVQA benchmark with full video clips. We also propose the task of multimodal entailment tree generation for the assessment of generated tree reasoning quality, establishing an information-theoretic evaluation method grounded in informal logic theory. Experimental results suggest that such interpretable, neuro-symbolic approaches to video understanding are a strong alternative to existing methods and present exciting directions for future research.

### v4

@@ 7 Conclusion
We introduce the first neuro-symbolic entailment tree generator for multimodal content to improve the robustness, reliability, and interpretability of video-language understanding systems. We propose the task of multimodal entailment tree generation for the assessment of generated tree reasoning quality, establishing an information-theoretic evaluation method grounded in informal logic theory. We show that our approach achieves state-of-the-art results on the zero-shot TVQA benchmark with full video clips, illustrating the potential for generated reasoning traces to improve downstream video-language understanding task performance. We show that interpretable, neuro-symbolic approaches to video understanding like TV-TREES are a strong alternative to existing methods, provide substantial new benefits, and highlight exciting directions for future research.
This paper is presented as an initial exploration into multimodal neuro-symbolic systems, and so there are many exciting avenues for future development and research. Individual components of this system could be improved for better performance on TVQA and related tasks - for instance, the vision querying system is fairly end-to-end, and semantically deconstructing the frames using a visual semantic role labeling model and using its outputs as evidence could result in a more sophisticated and transparent logical system. In future work, we also hope to explore the possibility of producing a collection of entailment trees pertaining to the same video clip and aggregating them to produce a comprehensive knowledge graph of the full video. We also hope to explore ways to improve computational efficiency and cost of using the system. Finally, there is significant room for future work in decomposing natural language text for entailment tree generation, and for establishing entailment between premises and grounded multimodal evidence.
@@ 8 Limitations
Acknowledgements
This work has been supported in part by the U.S. National Science Foundation under grant NSF 2204926.